OpenStudy (anonymous):

find the equqation of the line that passes through the point (4,-1) and (-2,3) Show all working please

OpenStudy (anonymous):

for the slope m=3+1/-2-4 =-2/3 but i dont know how i should plug in my points for the equation of the line y-y1=m(x-x1)

OpenStudy (anonymous):

i am suppose to plug in the actual numbers or what?

OpenStudy (anonymous):

y-(-1)=-2/3(x-4)

OpenStudy (anonymous):

so we could use any of the points and get the same answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh okay thank you thats what i didnt remeber